OpenStudy (anonymous):

Convert (126.14) from base 8 to decimal

OpenStudy (callisto):

1*8^2+2*8+6+1*8^(-1)+4*8^(-2) = 1379/16

OpenStudy (anonymous):

Explain!

OpenStudy (ash2326):

we have 1 2 6 . 1 4 2 1 0 . -1 -2 These are the places to convert to decimal \[1\times 8^2+ 2\times 8^1+6 \times 8^0+ 1\times 8^{-1} + 4 \times 8^{-2}\] we get \[64+16+6+1 \times 0.125+ 4 \times 0.015625\] we gat \[86.1875\]

OpenStudy (ash2326):

*get